zarasht:

A sample of argon has a volume of 1.2 L at STP. If the temperature is increased to 22°C and the pressure is lowered to 0.80 atm, what will the new volume be, in L?

justjm:

Use combined gas law \(\frac{P_{1}V_{1}}{T_{1}}=\frac{P_{2}V_{2}}{T_{2}}\) V1 = 1.2L STP means P1 = 1 ATM and T1 = 273 K Now you have to find V2 P2=0.80 atm and T = 273+22 K, *remember you must convert to kelvins*

zarasht:

Do i get the answer by dividing P2 by T2?

justjm:

\(\frac{1\cdot1.2}{273}=\frac{V_2\cdot0.8}{\left(22+273\right)}\) Solve using algebra (cross multiplication)

zarasht:

thanks